# An Agentic Framework with LLMs for Solving Complex Vehicle Routing Problems

### v1

@@ 1 Introduction
Vehicle routing problems (VRPs) are fundamental to industrial and commercial applications such as logistics (Bochtis & Sørensen, 2010; Konstantakopoulos et al., 2022) and transportation (Cattaruzza et al., 2017; Zhang et al., 2022), yet they remain challenging to solve due to their diverse variants with intricate real-world constraints. Traditional approaches (Furnon & Perron, ; Helsgaun, 2017; Vidal, 2022; Wouda et al., 2024) often require substantial expert effort, either to translate problem statements into mathematical formulations or to design specialized algorithms. Although recent neural solvers (Kool et al., 2018; Kwon et al., 2020) alleviate the dependence on domain knowledge, they still require a certain degree of manual adaptation to address more complex VRPs.
@@
More recently, large language models (LLMs) (Zhao et al., 2023), with their strong natural language understanding and code generation capabilities, may offer a promising avenue for automation, reducing reliance on manual effort and enabling flexible solver development across diverse VRP variants. Some early attempts (Yang et al., 2024; Liu et al., 2024b) directly prompt LLMs to generate solutions but fall short in terms of solution optimality and feasibility. Other approaches (Romera-Paredes et al., 2024) explore the use of LLMs to generate programming code as a proxy for addressing challenges in VRP optimization, which can be broadly categorized into two directions. The first direction centers on evolving basic heuristics tailored for conventional VRPs, with representative examples including EoH (Liu et al., 2024a) and ReEvo (Ye et al., 2024). In contrast, the second direction emphasizes developing general frameworks capable of handling diverse VRP variants, making it more practical and application-oriented.
@@
Early research efforts have started to address this challenging second direction. Their workflow generally comprises two phases: framework-design, in which the architecture and generation strategy are specified (see Section 3), and framework-execution, in which the resulting framework is deployed to solve diverse problem instances. ARS (Li et al., 2025a) constructs constraint-checking functions by retrieving and adapting templates from a predefined constraint library, while DRoC (Jiang et al., 2025) employs a retrieval-augmented generation (RAG) strategy to produce code that invokes OR-Tools (Furnon & Perron, ) for problem solving. Although both ARS and DRoC can handle complex VRPs, these module-level generation methods are not self-contained, depending on handcrafted code modules or external solvers during framework-design, and not fully automated, as they still require human involvement to extract instance-specific information during framework-execution. This dependence may introduce misalignment between LLM-generated code and external systems, which can result in execution errors and reduced solution feasibility. In contrast, SGE (Iklassov et al., 2024) achieves self-containment but is limited to relatively simple problems like the Traveling Salesman Problem (TSP), as it lacks effective mechanisms for handling complex constraints and fails to provide full automation or reliable code and solution validity. In this paper, as summarized in Table 1, we address these limitations by proposing a general framework of collaborative LLM-empowered agents that can tackle complex VRPs with self-containment, full automation, and high trustworthiness in both code and solutions.
@@
ARS
(Li et al., 2025a)
DRoC
(Jiang et al., 2025)
SGE
(Iklassov et al., 2024)
AFL
(This Work)
@@ 2 Preliminaries
The mapping between constraints and their corresponding VRPLIB fields is summarized in Table 2, with further details on each field provided in Table 9 in the Appendix. Our AFL directly takes VRPLIB-format instances as input.
@@ 3.4 Subtask 3: Solution Derivation
4 Experiment
We first evaluate AFL against traditional and neural approaches on 48 standard benchmarks incorporating common constraints such as capacity (C), duration limit (L), time window (TW), open route (O), multi depot (MD), backhaul (B), and mixture backhaul (MB), which are widely used to assess traditional algorithms. We then extend the evaluation to 8 practical electric (E) VRPs, which remain challenging for traditional solvers. Next, we benchmark AFL against LLM-based approaches, assessing code reliability, solution feasibility, and overall performance. We also conduct ablation studies to examine the effectiveness of our agentic design. Finally, we evaluate 4 additional open benchmarks, including TSP, ATSP, ACVRP, and SOP, to demonstrate the broad applicability of our framework. We run all traditional and LLM-based approaches on 16 CPU cores, with GPT-4.1 (OpenAI, 2024) as the default LLM used in AFL.
4.1 Comparison on Standard Benchmark
@@
HGS-PyVRP
CVRP
10.37 10.40 15.62 20.80
CVRPL
10.59 10.40 15.77 20.80
@@
HGS-PyVRP
CVRPTW
16.03 10.40 25.42 20.80
OCVRP
6.51 10.40 9.73 20.80
@@
HGS-PyVRP
CVRPLTW
16.36 10.40 25.76 20.80
OCVRPL
6.51 10.40 9.72 20.80
@@
HGS-PyVRP
OCVRPTW
10.51 10.40 16.93 20.80
OCVRPLTW
10.51 10.40 16.93 20.80
@@
HGS-PyVRP
CVRPB
9.69 10.40 14.38 20.80
CVRPBL
10.19 10.40 14.78 20.80
@@
HGS-PyVRP
CVRPBTW
18.29 10.40 29.47 20.80
OCVRPB
6.90 10.40 10.34 20.80
@@
HGS-PyVRP
CVRPBLTW
18.36 10.40 29.03 20.80
OCVRPBL
6.90 10.40 10.34 20.80
@@
HGS-PyVRP
OCVRPBTW
11.67 10.40 19.16 20.80
OCVRPBLTW
11.67 10.40 19.16 20.80
@@
We compare AFL with the traditional solvers HGS (implemented in PyVRP) (Vidal, 2022; Wouda et al., 2024) and OR-Tools (Furnon & Perron, ), as well as the neural solver RF-POMO (Berto et al., 2025b). The experimental settings and testing data follow Berto et al. (2025b), including 1,000 instances for each problem. Note that our objective is not to surpass SOTA solvers on conventional VRPs, which reflect decades of expert effort, but to develop a fully automated and self-contained framework for tackling complex VRPs. Therefore, in the comparison shown in Table 3, we regard a 3% relative gap with respect to SOTA solvers as an acceptable criterion. We report the results of AFL after 500, 2,000, and 10,000 iterations of solution improvement. The runtime of AFL is measured as the sum of the problem description and solution derivation phases. The code generation phase, analogous to the training phase of a neural solver, is excluded, since once the solver code is produced, it can be reused across instances without incurring repeated generation overhead. Results are partly shown in Table 3, with the rest provided in Table 10 in the Appendix due to space limitations.
4.2 Comparison on Practical Benchmark
Traditional solvers like HGS (Vidal, 2022; Wouda et al., 2024) and OR-Tools (Furnon & Perron, ) are inherently constrained by their internal implementations and cannot be directly adapted to new problem settings without substantial modifications to the core codebase, whereas AFL can naturally accommodate practical VRPs. To demonstrate this, we conduct experiments on a widely used benchmark for ECVRPTW (Schneider et al., 2014), a representative variant of the electric vehicle routing problem (EVRP). Specifically, the dataset contains 36 small instances with 5, 10, and 15 customers, as well as 56 large instances with 100 customers. To further assess generality, we extend this benchmark to 7 additional EVRP variants, namely ECVRP, ECVRPL, EOVRPL, EOCVRP, EOCVRPTW, ECVRPLTW, and EOCVRPLTW, enabling a more comprehensive evaluation across diverse and challenging problem settings. Given the intrinsic difficulty of these problems and the lack of directly applicable advanced solvers, we adopt ACO and Greedy as baselines, as both are widely recognized flexible heuristics for complex VRPs. For ACO, the number of improvement steps is fixed at 500. The results in Table 4 demonstrate the consistent effectiveness of AFL. Although ACO is executed with 500 improvement steps, our framework attains better objective values in shorter runtimes. These empirical findings highlight the superiority of AFL on complex and practical VRPs, where traditional solvers often face limitations.
@@ 4.3 Comparison with LLM-based Solver
We compare AFL in trustworthiness and performance with representative LLM-based approaches for diverse VRPs (above 16 variants plus TSP), namely SGE (Iklassov et al., 2024) and DRoC (Jiang et al., 2025), while excluding ARS (Li et al., 2025a) due to the unavailability of its source code.
@@
To assess performance, we further evaluate AFL on the problem classes (i.e., TSP, CVRP, and CVRPL) solvable by SGE and DRoC. Specifically, we conduct experiments on TSPLib (Reinelt, 1991), a real-world TSP benchmark containing 50 instances with sizes ranging from 50 to 1,000 customers, and on CVRPLib (Uchoa et al., 2017), a real-world CVRP benchmark containing 100 instances with sizes ranging from 100 to 1,000 customers. We further report the results of ReEvo (Ye et al., 2024) on TSPLib and CVRPLib, as both TSP and CVRP are solved in the original paper, to provide a broader assessment. For CVRPL, we adopt the same benchmark setting as in Section  3. The results are shown in Table 6, where gaps for TSPLib and CVRPLib are reported relative to their best-known solutions, while CVRPL gaps are measured against HGS. DRoC is able to solve CVRPLib instances only with fewer than 500 customers within the 10-hour time limit. Across all evaluated benchmarks, AFL consistently outperforms both SGE, DRoC, and ReEvo.
@@ 4.4 Comparison with Promoting Strategy
Table 7: Gap comparison across different prompting strategies.
CVRP VRPL VRPTW OVRP VRPLTW OVRPL OVRPTW OVRPLTW
50 100 50 100 50 100 50 100 50 100 50 100 50 100 50 100
Self-Verification 7.61% 9.47% 6.80% 8.64% 5.71% 9.25% 3.54% 5.96%
COT 8.39% 12.04% 7.37% 10.51% 56.84% 93.83% 62.06% 105.86%
@@ 4.6 Broad Applicability
We further evaluate AFL on four additional open benchmarks: TSP, ATSP, ACVRP, and SOP. Results for TSP are presented in Table 6. ATSP and ACVRP capture asymmetric routing, which frequently arises in real-world applications, while SOP introduces precedence-constrained path planning, another realistic and challenging setting. Specifically, the ATSP benchmark (Johnson & McGeoch, 1997) contains 18 instances ranging from 17 to 443 nodes. The ACVRP dataset (Helsgaun, 2017) provides 120 capacity-constrained cases with asymmetric distance matrices covering 16 to 200 customers. The SOP benchmark (Renaud et al., 1996) includes 39 precedence-constrained instances with sizes between 9 and 380 nodes. Experimental results for ATSP, ACVRP, and SOP are presented in Table 12, Table 13, and Table 14 in the Appendix, respectively. By achieving competitive performance across these diverse datasets, AFL demonstrates broad applicability, indicating that our framework has the potential to be extended to a wider range of problem variants.
@@ A.1 ML for VRPs
Existing neural approaches to solving VRPs generally follow two directions: learning to construct and learning to improve them. In the construction paradigm, a model either directly generates a feasible solution (Vinyals et al., 2015; Bello et al., 2017) or produces a probability heatmap (Joshi et al., 2021; Sun & Yang, 2023) from which a solution is sampled. AM (Kool et al., 2018) first introduces attention mechanisms for VRPs, and POMO (Kwon et al., 2020) subsequently exploits multiple optima to enhance solution quality, inspiring a series of extensions and refinements (Drakulic et al., 2023; Luo et al., 2023; Berto et al., 2025a). Recently, increasing attention has been devoted to improving generalization (Joshi et al., 2022; Zhou et al., 2023; Gao et al., 2024) and scalability (Li et al., 2021; Ouyang et al., 2025), addressing complex constraints (Bi et al., 2024; Li et al., 2025b), and handling multiple VRP variants within a single framework (Zhou et al., 2024; Berto et al., 2025b; Pan et al., 2025), though these efforts still fall short of practical deployment. In the improvement paradigm, an initial solution is iteratively refined using learned local-search (Wu et al., 2021; Ma et al., 2023) or destroy–repair strategies (Hottung & Tierney, 2022) to progressively reduce cost and enhance solution quality.
@@ A.2 LLM for VRPs
One line of work uses LLMs to directly generate or improve VRP solutions. For example, OPRO (Yang et al., 2024) attempts to construct solutions outright with an LLM, while LEMA (Liu et al., 2024b) leverages the LLM to perform the genetic search itself. However, they fall short in terms of solution quality and feasibility. Another line of research applies LLMs to generate code for VRPs, which can be broadly categorized into two directions: evolving basic heuristics for conventional VRPs and developing general frameworks for complex VRPs.
In evolving basic heuristics, LLMs are employed to iteratively evolve a simple or existing heuristic within fixed templates or established solvers for conventional VRPs. Early work such as EOH (Liu et al., 2024a) adopts a population-based framework with fixed templates for heuristic evolution. ReEvo (Ye et al., 2024) further combines evolutionary search with LLM reflections to provide verbal feedback and enhance search efficiency. More recently, HSEvo (Dat et al., 2025) and MCTS-AHD (Zheng et al., 2025) have advanced this line of research. However, these approaches remain tailored to specific problems, limiting their generality across VRP variants with practical constraints.
In developing general frameworks, LLMs are leveraged as knowledgeable developers to generate function modules with distinct roles, enabling the framework to address diverse and complex VRPs. In ARS (Li et al., 2025a) and DRoC (Jiang et al., 2025), LLMs generate specific functions to adapt frameworks to different variants, typically relying on handcrafted modules or external solvers. However, aligning LLM-generated code with these components remains challenging. In ARS, the integration between the LLM and the base algorithm is weak: the handcraft improvement operator and the constraint validation functions generated by the LLM are decoupled, and feasibility can only be verified after optimization, often leading to infeasible solutions. In DRoC, the LLM produces code that invokes OR-Tools without genuine knowledge of its internals, relying instead on external retrieval, which may undermine code reliability and reduce solution feasibility. In contrast, SGE (Iklassov et al., 2024) eliminates the need for predefined modules or solvers by directly generating code end-to-end. However, due to the inherent complexity of full code generation and the absence of effective constraint-handling mechanisms, its applicability is restricted to TSP. Moreover, it still depends on handcrafted extraction of instance information (e.g., inputs and constraints), and its overall performance remains uncompetitive.

### v2

@@ 1 Introduction
Vehicle routing problems (VRPs) are fundamental to industrial and commercial applications such as logistics (Bochtis and Sørensen, 2010; Konstantakopoulos et al., 2022) and transportation (Cattaruzza et al., 2017; Zhang et al., 2022), yet they remain challenging to solve due to their diverse variants with intricate real-world constraints. Traditional approaches (Furnon and Perron, 2025; Helsgaun, 2017; Vidal, 2022; Wouda et al., 2024) often require substantial expert effort, either to translate problem statements into mathematical formulations or to design specialized algorithms. Although recent neural solvers (Kool et al., 2018; Kwon et al., 2020) alleviate the dependence on domain knowledge, they still require a certain degree of manual adaptation to address more complex VRPs.
@@
More recently, large language models (LLMs) (Zhao et al., 2023), with their strong natural language understanding and code generation capabilities (Ma et al., 2026), may offer a promising avenue for automation, reducing reliance on manual effort and enabling flexible solver development across diverse VRP variants. Some early attempts (Yang et al., 2024; Liu et al., 2024b) directly prompt LLMs to generate solutions but fall short in terms of solution optimality and feasibility. Other approaches (Romera-Paredes et al., 2024) explore the use of LLMs to generate programming code as a proxy for addressing challenges in VRP optimization, which can be broadly categorized into two directions. The first direction centers on evolving basic heuristics tailored for conventional VRPs, with representative examples including EoH (Liu et al., 2024a) and ReEvo (Ye et al., 2024). In contrast, the second direction emphasizes developing general frameworks capable of handling diverse VRP variants, making it more practical and application-oriented.
@@
Early research efforts have started to address this challenging second direction. Their workflow generally comprises two phases: framework-design, in which the architecture and generation strategy are specified (see Section 3), and framework-execution, in which the resulting framework is deployed to solve diverse problem instances. ARS (Li et al., 2025a) constructs constraint-checking functions by retrieving and adapting templates from a predefined constraint library, while DRoC (Jiang et al., 2025b) employs a retrieval-augmented generation (RAG) strategy to produce code that invokes OR-Tools (Furnon and Perron, 2025) for problem solving. Although both ARS and DRoC can handle complex VRPs, these module-level generation methods are not self-contained, depending on handcrafted code modules or external solvers during framework-design, and not fully automated, as they still require human involvement to extract instance-specific information during framework-execution. This dependence may introduce misalignment between LLM-generated code and external systems, which can result in execution errors and reduced solution feasibility. In contrast, SGE (Iklassov et al., 2024) achieves self-containment but is limited to relatively simple problems like the Traveling Salesman Problem (TSP), as it lacks effective mechanisms for handling complex constraints and fails to provide full automation or reliable code and solution validity. In this paper, as summarized in Table 1, we address these limitations by proposing a general framework of collaborative LLM-empowered agents that can tackle complex VRPs with self-containment, full automation, and high trustworthiness in both code and solutions.
@@
ARS (Li et al., 2025a) DRoC (Jiang et al., 2025b) SGE (Iklassov et al., 2024) AFL (This Work)
@@ 2 Preliminaries
The mapping between constraints and their corresponding VRPLIB fields is summarized in Table 2, with further details on each field provided in Table 8 in the Appendix. Our AFL directly takes VRPLIB-format instances as input. We also evaluate AFL on JSON and CSV formats to demonstrate its robustness to different data representations. The results are reported in Section C.11 and Table 17.
@@ 3.4 Subtask 3: Solution Derivation
HGS-PyVRP CVRP 10.37 10.40 15.62 20.80 CVRPL 10.59 10.40 15.77 20.80
@@
HGS-PyVRP CVRPTW 16.03 10.40 25.42 20.80 OCVRP 6.51 10.40 9.73 20.80
@@
HGS-PyVRP CVRPLTW 16.36 10.40 25.76 20.80 OCVRPL 6.51 10.40 9.72 20.80
@@
HGS-PyVRP OCVRPTW 10.51 10.40 16.93 20.80 OCVRPLTW 10.51 10.40 16.93 20.80
@@
HGS-PyVRP CVRPB 9.69 10.40 14.38 20.80 CVRPBL 10.19 10.40 14.78 20.80
@@
HGS-PyVRP CVRPBTW 18.29 10.40 29.47 20.80 OCVRPB 6.90 10.40 10.34 20.80
@@
HGS-PyVRP CVRPBLTW 18.36 10.40 29.03 20.80 OCVRPBL 6.90 10.40 10.34 20.80
@@
HGS-PyVRP OCVRPBTW 11.67 10.40 19.16 20.80 OCVRPBLTW 11.67 10.40 19.16 20.80
@@
4 Experiment
We first evaluate AFL against traditional and neural approaches on 48 standard benchmarks incorporating common constraints such as capacity (C), duration limit (L), time window (TW), open route (O), multi depot (MD), backhaul (B), and mixture backhaul (MB), which are widely used to assess traditional algorithms. We then extend the evaluation to 8 practical electric (E) VRPs, which remain challenging for traditional solvers. Next, we benchmark AFL against LLM-based approaches, assessing code reliability, solution feasibility, and overall performance. We also conduct ablation studies to examine the effectiveness of our agentic design. In addition, we report comparisons under different prompt strategies in the main text. We further evaluate AFL’s robustness across different LLM backbones, and these results are presented in Section C.10 and Table 16 in the Appendix. Finally, we evaluate 4 additional open benchmarks, including TSP, ATSP, ACVRP, and SOP, to demonstrate the broad applicability of our framework. All experiments were conducted via the OpenAI API using GPT-4.1 (OpenAI, 2024) on a server equipped with an AMD EPYC 7702P CPU and 64 GB of RAM, without GPU acceleration. All results shown below are derived using the same heuristic for the same variant on a given instance. This setting is more practical for real-world applications and significantly reduces code-generation overhead. The code and dataset is available at https://github.com/ZHANG-NI/AFL.git
@@ 4 Experiment
4.1 Comparison on Standard Benchmark
We compare AFL with the traditional solvers HGS (implemented in PyVRP) (Vidal, 2022; Wouda et al., 2024) and OR-Tools (Furnon and Perron, 2025), as well as the neural solver RF-POMO (Berto et al., 2025b). The experimental settings and testing data follow Berto et al. (2025b), including 1,000 instances for each problem. Note that our objective is not to surpass SOTA solvers on conventional VRPs, which reflect decades of expert effort, but to develop a fully automated and self-contained framework for tackling complex VRPs. Therefore, in the comparison shown in Table 3, we regard a 3% relative gap with respect to SOTA solvers as an acceptable criterion. We report the results of AFL after 500, 2,000, and 10,000 iterations of solution improvement. The runtime of AFL is measured as the sum of the problem description and solution derivation phases. The code generation phase, analogous to the training phase of a neural solver, is excluded, since once the solver code is produced, it can be reused across instances without incurring repeated generation overhead, and the runtime analysis of code generation phase is shown in Section C.4 in Appendix. Results are partly shown in Table 3, with the rest provided in Table 11 in the Appendix due to space limitations.
4.2 Comparison on Practical Benchmark
Traditional solvers like HGS (Vidal, 2022; Wouda et al., 2024) and OR-Tools (Furnon and Perron, 2025) are inherently constrained by their internal implementations and cannot be directly adapted to new problem settings without substantial modifications to the core codebase, whereas AFL can naturally accommodate practical VRPs. To demonstrate this, we conduct experiments on a widely used benchmark for ECVRPTW (Schneider et al., 2014), a representative variant of the electric vehicle routing problem (EVRP). Specifically, the dataset contains 36 small instances with 5, 10, and 15 customers, as well as 56 large instances with 100 customers. To further assess generality, we extend this benchmark to 7 additional EVRP variants, namely ECVRP, ECVRPL, EOVRPL, EOCVRP, EOCVRPTW, ECVRPLTW, and EOCVRPLTW, enabling a more comprehensive evaluation across diverse and challenging problem settings. Given the intrinsic difficulty of these problems and the lack of directly applicable advanced solvers, we adopt ACO and Greedy as baselines, as both are widely recognized flexible heuristics for complex VRPs. For ACO, the number of improvement steps is fixed at 500. The results in Table 4 demonstrate the consistent effectiveness of AFL. Although ACO is executed with 500 improvement steps, our framework attains better objective values in shorter runtimes. These empirical findings highlight the superiority of AFL on complex and practical VRPs, where traditional solvers often face limitations.
@@ 4.3 Comparison with LLM-based Solver
We compare AFL in trustworthiness and performance with representative LLM-based approaches for diverse VRPs (above 16 variants plus TSP), namely SGE (Iklassov et al., 2024) and DRoC (Jiang et al., 2025b), while excluding ARS (Li et al., 2025a) due to the unavailability of its source code.
@@
To assess performance, we further evaluate AFL on the problem classes (i.e., TSP, CVRP, and CVRPL) solvable by SGE and DRoC. Specifically, we conduct experiments on TSPLib (Reinelt, 1991), a real-world TSP benchmark containing 50 instances with sizes ranging from 50 to 1,000 customers, and on CVRPLib (Uchoa et al., 2017), a real-world CVRP benchmark containing 100 instances with sizes ranging from 100 to 1,000 customers. We further report the results of ReEvo (Ye et al., 2024) on TSPLib and CVRPLib, as both TSP and CVRP are solved in the original paper, to provide a broader assessment. For CVRPL, we adopt the same benchmark setting as in Section  4.1. The results are shown in Table 6, where gaps for TSPLib and CVRPLib are reported relative to their best-known solutions, while CVRPL gaps are measured against HGS. DRoC is able to solve CVRPLib instances only with fewer than 500 customers within the 10-hour time limit. Across all evaluated benchmarks, AFL consistently outperforms both SGE, DRoC, and ReEvo.
@@ 4.6 Broad Applicability
We further evaluate AFL on four additional open benchmarks: TSP, ATSP, ACVRP, and SOP. Results for TSP are presented in Table 6. ATSP and ACVRP capture asymmetric routing, which frequently arises in real-world applications, while SOP introduces precedence-constrained path planning, another realistic and challenging setting. All three variants are formulated using non-Euclidean distance metrics. Specifically, the ATSP benchmark (Johnson and McGeoch, 1997) contains 18 instances ranging from 17 to 443 nodes. The ACVRP dataset (Helsgaun, 2017) provides 120 capacity-constrained cases with asymmetric distance matrices covering 16 to 200 customers. The SOP benchmark (Renaud et al., 1996) includes 39 precedence-constrained instances with sizes between 9 and 380 nodes. Experimental results for ATSP, ACVRP, and SOP are presented in Table 13, Table 14, and Table 15 in the Appendix, respectively. By achieving competitive performance across these diverse datasets, AFL demonstrates broad applicability, indicating that our framework has the potential to be extended to a wider range of problem variants.
@@ 5 Conclusion
Acknowledgments and Disclosure of Funding
This research is supported by the National Research Foundation, Singapore under the AI Singapore Programme (AISG Award No: AISG3-RPGV-2025-017).
@@ A.1 ML for VRPs
Existing neural approaches to solving VRPs generally follow two directions: learning to construct and learning to improve them. In the construction paradigm, a model either directly generates a feasible solution (Vinyals et al., 2015; Bello et al., 2017) or produces a probability heatmap (Joshi et al., 2021; Sun and Yang, 2023) from which a solution is sampled. AM (Kool et al., 2018) first introduces attention mechanisms for VRPs, and POMO (Kwon et al., 2020) subsequently exploits multiple optima to enhance solution quality, inspiring a series of extensions and refinements (Drakulic et al., 2023; Luo et al., 2023; Berto et al., 2025a; Zhang and Cao, 2025). Recently, increasing attention has been devoted to improving generalization (Joshi et al., 2022; Zhou et al., 2023; Gao et al., 2024) and scalability (Li et al., 2021; Ouyang et al., 2025), addressing complex constraints (Bi et al., 2024; Li et al., 2025b), and handling multiple VRP variants within a single framework (Zhou et al., 2024; Berto et al., 2025b; Pan et al., 2025), though these efforts still fall short of practical deployment. In the improvement paradigm, an initial solution is iteratively refined using learned local-search (Wu et al., 2021; Ma et al., 2023) or destroy–repair strategies (Hottung and Tierney, 2022) to progressively reduce cost and enhance solution quality.
@@ A.2 LLM for VRPs
One line of work uses LLMs to directly generate or improve VRP solutions. For example, OPRO (Yang et al., 2024) attempts to construct solutions outright with an LLM, LEMA (Liu et al., 2024b) leverages the LLM to perform the genetic search itself, and Jiang et al. (2025a) finetune LLM to construct in the end to end manner. However, they fall short in terms of solution quality and feasibility. Another line of research applies LLMs to generate code for VRPs, which can be broadly categorized into two directions: evolving basic heuristics for conventional VRPs and developing general frameworks for complex VRPs.
In evolving basic heuristics, LLMs are employed to iteratively evolve a simple or existing heuristic within fixed templates or established solvers for conventional VRPs. Early work such as EOH (Liu et al., 2024a) adopts a population-based framework with fixed templates for heuristic evolution. ReEvo (Ye et al., 2024) further combines evolutionary search with LLM reflections to provide verbal feedback and enhance search efficiency. More recently, HSEvo (Dat et al., 2025) and MCTS-AHD (Zheng et al., 2025) have advanced this line of research. Zhu et al. (2026) uses evolution framework to bridge synthetic and real routing problems. However, these approaches remain tailored to specific problems, limiting their generality across VRP variants with practical constraints.
In developing general frameworks, LLMs are leveraged as knowledgeable developers to generate function modules with distinct roles, enabling the framework to address diverse and complex VRPs. In ARS (Li et al., 2025a) and DRoC (Jiang et al., 2025b), LLMs generate specific functions to adapt frameworks to different variants, typically relying on handcrafted modules or external solvers. However, aligning LLM-generated code with these components remains challenging. In ARS, the integration between the LLM and the base algorithm is weak: the handcraft improvement operator and the constraint validation functions generated by the LLM are decoupled, and feasibility can only be verified after optimization, often leading to infeasible solutions. In DRoC, the LLM produces code that invokes OR-Tools without genuine knowledge of its internals, relying instead on external retrieval, which may undermine code reliability and reduce solution feasibility. In contrast, SGE (Iklassov et al., 2024) eliminates the need for predefined modules or solvers by directly generating code end-to-end. However, due to the inherent complexity of full code generation and the absence of effective constraint-handling mechanisms, its applicability is restricted to TSP. Moreover, it still depends on handcrafted extraction of instance information (e.g., inputs and constraints), and its overall performance remains uncompetitive.